OpenStudy (anonymous):

Simplify x^2-6x+9/x^2-9

OpenStudy (lgbasallote):

\[\frac{x^2 - 6x+9}{x^2 - 9} \implies \frac{(x-3)^2}{(x+3)(x-3)}\] does that help?

OpenStudy (anonymous):

so would the answer be \[x+6/x-6\]

OpenStudy (lgbasallote):

where did you get that?

OpenStudy (anonymous):

i guessed... it was one of my answers. i thought it made sense in my head

OpenStudy (lgbasallote):

look at my explanation it will make more sense

OpenStudy (lgbasallote):

\[\frac{(x-3)^2}{(x+3)(x-3)} \implies \frac{(x-3)(x-3)}{(x+3)(x-3)}\] do you see it now?

OpenStudy (anonymous):

so \[x-6/x-3\]

OpenStudy (anonymous):

how do you make the line for the fraction?

OpenStudy (lgbasallote):

stop guessing...you're not gonna make me say yes to the right answer lol...just observe...and think of cancellation

OpenStudy (anonymous):

sorry this test is starting to get to me. been taking it for 5 hours. but it'd be \[x-3/x+3\] I'm so tired i didn't even see you could easily cancel it out...